# Bootstrapped Control Limits for Score-Based Concept Drift Control Charts

### v1

@@ 3 Nested Bootstrap Procedure for Computing the CL
Comparing (3.17) and (3.18), we see that $\operatorname{Cov}_{0}[\mathbf{z}_{i}^{\,b,j}]=\operatorname{Cov}_{0}[\mathbf{z}_{i}]+[1-(1-\lambda)^{i}]^{2}\mathbf{I}(\boldsymbol{\theta}^{0})/{0.368n}$, where, from equations  (3.9) and (3.11), the term $I(\boldsymbol{\theta}^{0})/(0.368\,n)$ in the difference is precisely $\mathbb{E}_{0}[\operatorname{Cov}_{0}[\overline{\mathbf{s}}_{\mathrm{OOB}}^{\,b}\mid\hat{\boldsymbol{\theta}}^{\,b}]]$. Thus, this difference accounts for the variability in the mean $\overline{\mathbf{s}}_{\mathrm{OOB}}^{\,b}$ of the “population” $\mathcal{S}_{\mathrm{OOB}}^{\,b}$ (whose cardinality is $0.368\,n$) from which the inner bootstrap samples are drawn in Algorithm 1. This constitutes a nuanced instantiation of the $0.632$ bootstrap rule that is crucial for obtaining a correct CL. When this correction is omitted (in which case the scaling factor $k(\lambda,i,n)$ in (3.20) is replaced by $1$), we have observed empirically that the bootstrapped CL is typically far too large and the detection power is unnecessarily compromised.
@@
Several implementation aspects of the algorithm warrant further discussion. When fitting models to bootstrap samples in Step 3-II, it is important to use the same tuning parameters $\gamma$ that were selected in Step 1 rather than performing new CV, as this maintains consistency in the model structure across replicates. If the covariance matrices $\widehat{\boldsymbol{\Sigma}}$ in Step 2 and $\widehat{\boldsymbol{\Sigma}}^{b}$ in Step 3-III are poorly conditioned (e.g., when $\hat{\boldsymbol{\theta}}^{\,b}$ is high-dimensional and/or has highly correlated components), then $\epsilon\mathbf{I}$ for some small scalar $\epsilon$ should be added to them before they are inverted in (3.20) and (3.21). If this is done, it should be done consistently in Steps 2 and 3-III using the same value of $\epsilon$. Note that $\widehat{\boldsymbol{\Sigma}}^{b}$ will typically be more poorly conditioned than $\widehat{\boldsymbol{\Sigma}}$, since there are fewer distinct score vectors in $\mathbf{S}_{\mathcal{D}}^{b}$ than in $\mathbf{S}_{\mathcal{D}}$. Thus, whether $\widehat{\boldsymbol{\Sigma}}$ is replaced by $\widehat{\boldsymbol{\Sigma}}+\epsilon\mathbf{I}$ in Step 2 should depend on how poorly conditioned the $\widehat{\boldsymbol{\Sigma}}^{b}$ are in step 3-III. The value for the desired false-alarm probability $\alpha$ (e.g., 0.01, 0.001, etc) should be chosen to balance detection power (which decreases as $\alpha$ decreases) with false-alarm control (which decreases as $\alpha$ decreases). The number of bootstrap replicates $B_{O}$ and $B_{I}$ should be chosen with computational considerations in mind. Because the dominant cost comes from refitting the model to each outer bootstrap sample $\mathcal{D}^{b}$, we use a smaller $B_{O}$ together with a larger $B_{I}$ to obtain accurate quantile estimates at reasonable runtime. In all of our numerical simulations in Section 4 we used $B_{O}=100$, $B_{I}=200$, and set the false-alarm probability to $\alpha=0.001$.
@@
Concept drift is detected when $T_{n+i}$ exceeds its corresponding CL $CL_{i}$. The choice of the MEWMA smoothing parameter $\lambda$ also affects the monitoring sensitivity, and the tradeoff is the same as in any MEWMA monitoring procedure: Smaller values of $\lambda$ provide more powerful eventual detection of large shifts but can delay the detection of large shifts, whereas larger values provide quicker detection of large shifts but may fail to detect smaller shifts. Finally, it is worth mentioning that our Algorithm 1 exhibits a high degree of parallelism, in that the outer procedures for $i=1,2,\cdots,B_{O}$ are independent, allowing for straightforward parallelization. This is especially important, since the primary computational expense of Algorithm 1 is in fitting the model to each outer bootstrap sample $\mathcal{D}^{b}$.
@@ 4.1 Mixed Linear Population
where predictor values $x_{i}$ are uniformly sampled from $[-\sqrt{3},\sqrt{3}]$. The training data of size $n=2,000$ are generated exclusively from model (4.1), yielding an pre-shift sample governed by a single linear relationship. The future data to be monitored comprise 1,000 observations: the first 200 points (pre-shift) follow the same single linear model (4.1), while the remaining 800 points (post-shift) are generated from a mixture that draws $y_{i}$ from either (4.1) or (4.2) with equal probability 0.5, representing a shift in $\mathbb{P}(Y|\mathbf{X})$.
@@
We used MEWMA parameter $\lambda=0.01$, which we chose via 5-fold CV, and desired pointwise false-alarm probability $\alpha=0.001$. In Algorithm 1, we used $B_{O}=100$, $B_{I}=200$, and $\lambda=0.01$.
@@
Figure 1: Visualization of predictive relationships for the linear example before and after the shift. The blue points are a scatter plot of $y$ vs. $x$ for the pre-shift single linear model (4.1), and the orange points correspond to the post-shift mixture model  (4.1) and (4.2). The shift from a single component to a mixture alters $\mathbb{P}(Y|\mathbf{X})$, representing a structural shift in the predictive relationship that is relatively small but can still be detected by our approach.
@@
Figure 2: Typical monitoring results using our bootstrap CL approach for the linear mixture example showing the detection of concept drift at new observation number 258 (the shift first occurred at observation number 201). The $T^{2}$ statistics ($T_{i}$ from Eq. (3.21), represented by the blue line) remain below the control limit (dashed red line) prior to the shift at observation 201, and sharply exceed it shortly after the shift.
Figure 3: Comparison of the pointwise false-alarm rate for our bootstrap CL (blue curve) versus the CL of [20] (orange curve) averaged over 50 database replicates under the pre-shift predictive relationship (4.1). Our bootstrap CL much more accurately maintains the empirical false-alarm rate close to the desired $\alpha=0.001$.
@@
In Figure 1, we visualize the transition from the pre-shift predictive relationship (4.1) to the post-shift relationship that is a mixture of (4.1) and (4.2). Although the shift appears quite small in Figure 1, Figure 2 demonstrates that it is still detectable. Figure 2 shows a typical evolution of our $T^{2}$ monitoring statistic over 1,000 observations with the first 200 following the pre-shift distribution and the last 800 the post-shift distribution. The $T^{2}$ statistic remains below the CL prior to the shift and then begins to increase immediately following the shift at time 201 until it first exceeds the CL at time 258, demonstrating the method’s ability to detect the shift. This behaviour aligns with our derivation in Section 3, where changes in $\mathbb{P}(Y|\mathbf{X})$ were shown to induce non-zero means in the score vectors.
@@
To better assess false-alarm control, the blue curve in Figure 3 shows the pointwise false-alarm rate over the first 1,000 observations when no shift occurs, estimated as the average pointwise false-alarm rate over 50 replicates. On each replicate, a different training data set was generated, the ridge regression model was fit to these data, Algorithm 1 was used to compute the CL as a function of time, a new set of 1,000 observations was generated from the same distribution (representing no shift), and then the $T^{2}$ statistic was computed for each new observation and compared to its CL to determine if there was a signal at that time. Our method consistently maintains the false-alarm rate at close to the desired value $\alpha=0.001$. The orange curve shows the corresponding results from the baseline method of [20], which exhibits substantially inflated false-alarm rates exceeding 0.06, highlighting the more accurate false-alarm rate provided by our bootstrap CL.
@@
The discrepancy arises for two reasons. First, [20] used a constant CL, because without bootstrapping there is no clear way to compute a time-varying CL. Second, [20] fit their model on only the first half of the training data and compute the constant CL from the second half. The resulting CL is based on a much smaller effective sample size than in our nested bootstrapping procedure and cannot reliably provide correct Type I error control unless the training sample size is quite large. In contrast, our approach provides accurate false-alarm rate control even with the relatively small training sample size of 2,000 in this example.
@@
It should be noted that the intended application scenarios in [20] involved much larger training sample sizes and/or larger values for the desired $\alpha$. For $\alpha=0.01$ (for example), the method of [20] provides much more accurate false-alarm control than in this example with $\alpha=0.001$. With $\alpha=0.001$, the method of [20] generally requires that the size of the training subset used to compute the CL is at least 10,000 to accurately estimate the upper 0.001 quantile of the empirical distribution of the $T^{2}$ statistic. In this regard, our bootstrap CL approach can be viewed as an extension of the approach of [20] to smaller training sample sizes and/or larger desired $\alpha$ values.
@@ 4.2 Nonlinear Oscillator System
To demonstrate our method’s effectiveness in detecting changes in complex nonlinear predictive relationships, we consider a physics-based system that represents a variant of the two-degree-of-freedom nonlinear oscillator with a finite-extensibility coupling spring proposed by [7]. The system consists of two masses moving in one dimension, connected by nonlinear springs. Let $p_{1}(t)$ and $p_{2}(t)$ denote the positions at (continuous) time $t$ of the two masses having masses $m_{1}$ and $m_{2}$, respectively, and let $v_{1}(t)=\dot{p}_{1}(t)$ and $v_{2}(t)=\dot{p}_{2}(t)$ denote their velocities. The system evolves according to the following differential equations
$\ddot{p}_{1}(t)$ $=\frac{1}{m_{1}}(-k_{1}p_{1}(t)-c_{1}\dot{p}_{1}(t)+k_{3}\phi(p_{1}(t),p_{2}(t))),$ (4.4)
$\ddot{p}_{2}(t)$ $=\frac{1}{m_{2}}(-k_{2}p_{2}(t)-c_{2}\dot{p}_{2}(t)-k_{3}\phi(p_{1}(t),p_{2}(t))),$
where $\phi(p_{1},p_{2})=(p_{1}-p_{2})/(1+|p_{1}-p_{2}|)$ represents the nonlinear coupling force. We sample the oscillator uniformly on $[0,30]$ to build a training set $\{(\mathbf{X}_{i},y_{i})\}_{i=1}^{n}$ with observations of $n=3{,}000$. The feature vector $\mathbf{X}_{i}\in\mathbb{R}^{4}$ from the state of the system is
$\mathbf{X}_{i}=[p_{1}(t_{i}),v_{1}(t_{i}),p_{2}(t_{i}),v_{2}(t_{i})]^{\top},$ (4.5)
@@
$y_{i}=\frac{1}{2}(m_{1}v_{1}^{2}(t_{i})+m_{2}v_{2}^{2}(t_{i}))+\frac{1}{2}(k_{1}p_{1}^{2}(t_{i})+k_{2}p_{2}^{2}(t_{i}))+k_{3}\phi\left(p_{1}(t_{i}),p_{2}(t_{i})\right)+\epsilon_{i},$ (4.6)
where $\varepsilon_{i}\sim\mathcal{NID}(0,\sigma^{2})$ represents measurement noise. For the training data, we used the system parameters $m_{1}=1.0$, $m_{2}=2.0$, $k_{1}=1.0$, $k_{2}=2.0$, $k_{3}=1.5$, $c_{1}=0.1$, and $c_{2}=0.2$. This setting presents a more challenging monitoring problem than the previous example because the predictive relationship is inherently nonlinear, and the predictive model that we fit to capture this is a neural network (described below).
To simulate concept drift, we generate new observations from the same system but with modified physical parameters
$m_{1}^{\prime}=1.1m_{1},\quad m_{2}^{\prime}=1.2m_{2},\quad k_{1}^{\prime}=1.3k_{1},$ (4.7)
while holding the remaining parameters fixed. These changes represent interpretable real-world perturbations, such as increased mass (e.g., due to material accumulation) or stiffened springs (e.g., due to thermal effects). They alter both the system’s natural frequencies and the energy landscape, inducing subtle but systematic shifts in the predictive relationship that our monitoring framework aims to detect.
We fit a feed-forward neural network with one hidden layer of 4 ReLU units to model the predictive relationship. Let $\boldsymbol{\Theta}=(\mathbf{W}_{1},\boldsymbol{b}_{1},\boldsymbol{w},b)$, where $\mathbf{W}_{1}\in\mathbb{R}^{4\times 4}$ and $\boldsymbol{b}_{1}\in\mathbb{R}^{4}$ denote the weights and biases for the hidden layer, and let $\boldsymbol{w}\in\mathbb{R}^{4}$ and $b\in\mathbb{R}$ denote the weights and biases of the output layer. The resulting predictor is
$g_{\gamma}(\boldsymbol{\Theta};\mathbf{x})=\boldsymbol{w}^{\!\top}\sigma\!\bigl{(}\mathbf{W}_{1}\mathbf{x}+\boldsymbol{b}_{1}\bigr{)}+b,\qquad\sigma(\mathbf{z})=\max\{\mathbf{0},\mathbf{z}\},$
where the vector max operation is element-by-element and given training data, the parameter vector monitored for drift is ${\boldsymbol{\theta}}=({\boldsymbol{w}},{b})$. For large neural network models, [20] recommends defining ${\boldsymbol{\theta}}$ as only the parameters of the last layer to keep its dimension reasonable and since changes in the predictive relationship should be reflected as changes in the parameters of the last layer when the parameters of the earlier layers are fixed, and we adopt this convention here. When fitting the model to the training data the complete set of parameters are estimated, but for the purpose of monitoring the reduced dimension ${\boldsymbol{\theta}}$ with ${\mathbf{W}}_{1}$ and ${\boldsymbol{b}}_{1}$ viewed as fixed, the penalized log likelihood is
@@
$\ell_{\mathrm{pen}}({\boldsymbol{\theta}};\mathbf{x}_{i},y_{i})=-\frac{1}{2\sigma^{2}}\bigl{(}y_{i}-g_{\gamma}({\boldsymbol{\theta}};\mathbf{x}_{i})\bigr{)}^{2}-\frac{1}{2}\log\!\bigl{(}2\pi\sigma^{2}\bigr{)}-\frac{\gamma}{2n}\bigl{(}\|{\boldsymbol{w}}\|^{2}+b^{2}\bigr{)},$
with regularization parameter $\gamma=10^{-1}$. For an observation $(\mathbf{x}_{i},y_{i})$, the corresponding score vector is the gradient of this log-likelihood with respect to $\boldsymbol{\theta}$, evaluated at the fitted parameters $\hat{\boldsymbol{\theta}}=(\hat{\boldsymbol{w}},\hat{b})$, is
$\mathbf{s}(\hat{\boldsymbol{\theta}};\mathbf{x}_{i},y_{i})=\frac{y_{i}-g_{\gamma}(\hat{\boldsymbol{\theta}};\mathbf{x}_{i})}{\sigma^{2}}\begin{bmatrix}\sigma\!\bigl{(}\hat{\mathbf{W}}_{1}\mathbf{x}_{i}+\hat{\boldsymbol{b}}_{1}\bigr{)}\\[2.0pt] 1\end{bmatrix}-\frac{\gamma}{n}\begin{bmatrix}\hat{\boldsymbol{w}}\\[2.0pt] \hat{b}\end{bmatrix}\;\in\mathbb{R}^{5}.$
The explicit form is shown for clarity and the implementation does not hand-code this gradient. Instead, for each observation we call loss.backward() in PyTorch, letting automatic differentiation compute the score vector directly from the penalized loss. This keeps the monitoring code concise and readily adaptable to more complex network architectures.
After fitting the network on this baseline data, we run Algorithm 1 with $B_{O}=100$, $B_{I}=200$, $\alpha=0.001$, and $\lambda=0.01$. The subsequent monitoring stream contains $M=1{,}000$ observations, of which the final 800 incorporate the prescribed parameter shift (4.7).
(a) Low noise ($\sigma=0.03$, 5-fold CV $R^{2}\!\approx\!0.928$)
(b) High noise ($\sigma=0.3$, 5-fold CV $R^{2}\!\approx\!0.108$)

### v2

@@ 3.2 Stabilization and Detectability of the Control Limit
Theorem 3.3 establishes geometric stabilization for the population bootstrap quantile $q_{i,\alpha}(\mathcal{D})$. In practice, Algorithm 1 computes the control limit $\mathrm{CL}_{i}$ as an empirical $(1-\alpha)$ quantile based on $B_{O}\times B_{I}$ bootstrap replicates, which introduces additional Monte Carlo variability. Because Phase I uncertainty is captured by the outer bootstrap loop, the effective number of independent outer draws is $B_{O}$, and the Monte Carlo error is therefore typically dominated by an $O_{p}(B_{O}^{-1/2})$ term. The numerical results in Section 4, obtained with $B_{O}=100$ and $B_{I}=200$, indicate that this Monte Carlo error is negligible for practical purposes.
Remark 3.3.
@@
Equation (3.4) implies that $\operatorname{Cov}_{0}[\mathbf{z}_{i}^{\,b,j}]=\operatorname{Cov}_{0}[\mathbf{z}_{i}]+[1-(1-\lambda)^{i}]^{2}\mathbf{I}(\boldsymbol{\theta}^{0})/{0.368n}$, where, from equation  (A.8), the term $\mathbf{I}(\boldsymbol{\theta}^{0})/0.368n$ ($\mathbf{I}(\boldsymbol{\theta}^{0})$ is the Fisher information matrix) in the difference is precisely $\mathbb{E}_{0}[\operatorname{Cov}_{0}[\overline{\mathbf{s}}_{\mathrm{OOB}}^{\,b}\mid\hat{\boldsymbol{\theta}}^{\,b}]]$. Thus, this difference accounts for the variability in the mean $\overline{\mathbf{s}}_{\mathrm{OOB}}^{\,b}$ of the “population” $\mathcal{S}_{\mathrm{OOB}}^{\,b}$ (whose cardinality is 0.368n0.368\,n) from which the inner bootstrap samples are drawn in Algorithm 1. This constitutes a nuanced instantiation of the $0.632$ bootstrap rule that is crucial for obtaining a correct CL. When this correction is omitted (in which case the scaling factor $k(\lambda,i,n)$ in (3.5) is replaced by $1$), we have observed empirically that the bootstrapped CL is typically far too large and the detection power is unnecessarily compromised.
The preceding results establish accurate control limit calibration for the proposed score-based MEWMA chart under correct model specification. We now investigate whether the score-based monitoring mechanism continues to detect changes in the predictive relationship when the supervised learning model is misspecified. In particular, we focus on whether a distributional shift induces a nonzero mean in the monitored score vector, which is the fundamental signal driving detection. Let $\mathbb{P}_{0}$ denote the in-control joint distribution of $(\mathbf{X},Y)$, and let $p_{\boldsymbol{\theta}}(y\mid\mathbf{x})$ denote the working conditional density (possibly misspecified). Define the pseudo-true parameter $\boldsymbol{\theta}^{\star}:=\arg\max_{\boldsymbol{\theta}\in\Theta}\mathbb{E}_{\mathbb{P}_{0}}\!\left[\log p_{\boldsymbol{\theta}}(Y\mid\mathbf{X})\right],$ and let $\mathbf{s}(\boldsymbol{\theta};\mathbf{x},y):=\nabla_{\boldsymbol{\theta}}\log p_{\boldsymbol{\theta}}(y\mid\mathbf{x})$ denote the corresponding score vector. For a post-change distribution $\mathbb{P}_{1}\neq\mathbb{P}_{0}$, define $\boldsymbol{\theta}_{1}^{\star}$ analogously.
Theorem 3.4 (Detectability under model misspecification).
Assume that the pseudo-true parameters $\boldsymbol{\theta}^{\star}$ and $\boldsymbol{\theta}_{1}^{\star}$ are uniquely defined and satisfy the regularity conditions of Assumption A.3. Then
  1. (a)
    In-control score mean $\mathbb{E}_{\mathbb{P}_{0}}\!\left[\mathbf{s}(\boldsymbol{\theta}^{\star};\mathbf{X},Y)\right]=\mathbf{0}.$
  2. (b)
    If $\boldsymbol{\theta}_{1}^{\star}\neq\boldsymbol{\theta}^{\star}$, then $\mathbb{E}_{\mathbb{P}_{1}}\!\left[\mathbf{s}(\boldsymbol{\theta}^{\star};\mathbf{X},Y)\right]\neq\mathbf{0}.$
  3. (c)
    Under $\mathbb{P}_{1}$, the MEWMA statistic constructed from $\mathbf{s}(\boldsymbol{\theta}^{\star};\mathbf{X},Y)$ converges geometrically in mean to a nonzero vector.
Theorem 3.4 shows that detectability of concept drift does not rely on correct model specification. That is, any shift that changes the pseudo-true parameter produces a nonzero mean in the monitored score vector, yielding a detectable signal in the MEWMA statistic.
3.3 Practical Considerations and Computational Aspects
@@
Several implementation aspects of the algorithm warrant further discussion. When fitting models to bootstrap samples in Step 3-II, it is important to use the same tuning parameters $\gamma$ that were selected in Step 1 rather than performing new CV, as this maintains consistency in the model structure across replicates. If the covariance matrices $\widehat{\boldsymbol{\Sigma}}$ in Step 2 and $\widehat{\boldsymbol{\Sigma}}^{b}$ in Step 3-III are poorly conditioned (e.g., when $\hat{\boldsymbol{\theta}}^{\,b}$ is high-dimensional and/or has highly correlated components), then $\epsilon\mathbf{I}$ for some small scalar $\epsilon$ should be added to them before they are inverted in (3.5) and (3.6). If this is done, it should be done consistently in Steps 2 and 3-III using the same value of $\epsilon$. Note that $\widehat{\boldsymbol{\Sigma}}^{b}$ will typically be more poorly conditioned than $\widehat{\boldsymbol{\Sigma}}$, since there are fewer distinct score vectors in $\mathbf{S}_{\mathcal{D}}^{b}$ than in $\mathbf{S}_{\mathcal{D}}$. Thus, whether $\widehat{\boldsymbol{\Sigma}}$ is replaced by $\widehat{\boldsymbol{\Sigma}}+\epsilon\mathbf{I}$ in Step 2 should depend on how poorly conditioned the $\widehat{\boldsymbol{\Sigma}}^{b}$ are in step 3-III.
The choice of monitoring parameters $\alpha$ and $\lambda$ follows standard MEWMA practice. $\alpha$ is determined by the acceptable false-alarm rate, and $\lambda$ balances the trade-off between detecting small, gradual shifts versus large, abrupt shifts (Montgomery 2020). For the bootstrap parameters, $B_{O}$ governs the accuracy of estimating the variability due to model fitting, while $B_{I}$ captures the variability of future data. Since $B_{O}$ determines the number of model refits, it represents the primary computational constraint. However, we note that the $B_{O}$ replicates are independent and can be efficiently performed simultaneously across multiple cores or GPUs. For deep learning models (like the one in Section 4.2), the refitting process can leverage GPU acceleration to significantly reduce runtime. For $B_{I}$, the inner-loop computations are relatively inexpensive. Accordingly, we recommend that practitioners choose $B_{I}$ to be conservatively large (e.g., $B_{I}\geq 200$) and focus computational resources on maximizing $B_{O}$, which governs the dominant cost of model refitting.
Remark 3.4.
In some practical applications, new data may arrive in batches rather than as single observations. In such cases, we recommend randomly ordering the observations within the batch and processing them sequentially through the MEWMA recursion. Because the control limit $CL_{i}$ is a function of the time index $i$ relative to the start of monitoring (and the training data) but does not depend on the specific values of the new observations, the random ordering of a batch does not alter the threshold sequence.
@@
Concept drift is detected when $T_{n+i}$ exceeds its corresponding CL $CL_{i}$. The choice of the MEWMA smoothing parameter $\lambda$ also affects the monitoring sensitivity, and the tradeoff is the same as in any MEWMA monitoring procedure: Smaller values of $\lambda$ provide more powerful eventual detection of large shifts but can delay the detection of large shifts, whereas larger values provide quicker detection of large shifts but may fail to detect smaller shifts.
@@ 4.1 Mixed Linear Population
where predictor values $x_{i}$ are uniformly sampled from $\sqrt{3},\sqrt{3}]$. The training data of size $n=2,000$ are generated exclusively from model (4.1), yielding an pre-shift sample governed by a single linear relationship. The future data to be monitored comprise 1,000 observations: the first 200 points (pre-shift) follow the same single linear model (4.1), while the remaining 800 points (post-shift) are generated from a mixture that draws $y_{i}$ from either (4.1) or (4.2) with equal probability 0.5, representing a shift in $\mathbb{P}(Y|\mathbf{X})$.
@@
We used MEWMA parameter $\lambda=0.01$, which we chose via 5-fold CV, and desired pointwise false-alarm probability $\alpha=0.001$. In Algorithm 1, we used $B_{O}=100$ and $B_{I}=200$.
@@
Figure 4.1: Visualization of predictive relationships for the linear example before and after the shift. The blue points are a scatter plot of $y$ vs. $x$ for the pre-shift single linear model (4.1), and the orange points correspond to the post-shift mixture model  (4.1) and (4.2). The shift from a single component to a mixture alters $\mathbb{P}(Y|\mathbf{X})$, representing a structural shift in the predictive relationship that is relatively small but can still be detected by our approach.
(a) (b)
@@
Figure 4.2: (a) Typical monitoring results using our bootstrap CL approach for the linear mixture example showing the detection of concept drift at new observation number 258 (the shift first occurred at observation number 201). The $T^{2}$ statistics ($T_{i}$ from Eq. (3.6), represented by the blue line) remain below the control limit (dashed orange line) prior to the shift at observation 201, and sharply exceed it shortly after the shift. (b) Comparison of the pointwise false-alarm rate for our bootstrap CL (blue curve) versus the CL of Zhang et al. (2023) (orange curve) averaged over 50 database replicates under the pre-shift predictive relationship (4.1). Our bootstrap CL much more accurately maintains the empirical false-alarm rate close to the desired $\alpha=0.001$.
@@
In Figure 4.1, we visualize the transition from the pre-shift predictive relationship (4.1) to the post-shift relationship that is a mixture of (4.1) and (4.2). Although the shift appears quite small in Figure 4.1, Figure 4.2(a) demonstrates that it is still detectable. Figure 4.2(a) shows a typical evolution of our $T^{2}$ monitoring statistic over 1,000 observations, with the first 200 following the pre-shift distribution and the last 800 the post-shift distribution. The $T^{2}$ statistic remains below the CL prior to the shift and then begins to increase immediately following the shift at time 201 until it first exceeds the CL at time 258, demonstrating the method’s ability to detect the shift. This behaviour aligns with our derivation in Section 3, where changes in $\mathbb{P}(Y|\mathbf{X})$ were shown to induce non-zero means in the score vectors.
@@
To better assess the empirical false-alarm control, the blue curve in Figure 4.2(b) shows the pointwise false-alarm rate over the first 1,000 observations when no shift occurs, estimated as the average pointwise false-alarm rate over 50 replicates. On each replicate, a different training data set was generated, the ridge regression model was fit to these data, Algorithm 1 was used to compute the CL as a function of time, a new set of 1,000 observations was generated from the same distribution (representing no shift), and then the $T^{2}$ statistic was computed for each new observation and compared to its CL to determine if there was a signal at that time. Our method consistently maintains the false-alarm rate at close to the desired value $\alpha=0.001$. The orange curve shows the corresponding results from the baseline method of Zhang et al. (2023), which exhibits substantially inflated FAR exceeding 0.06, highlighting the more accurate false-alarm rate provided by bootstrap CL.
@@
The discrepancy arises for two reasons. First, Zhang et al. (2023) used a constant CL, because without bootstrapping there is no clear way to compute a time-varying CL. Second, Zhang et al. (2023) fit their model on only the first half of the training data and compute the constant CL from the second half. The resulting CL is based on a much smaller effective sample size than in our nested bootstrapping procedure and cannot reliably provide correct Type I error control unless the training sample size is quite large. In contrast, our approach provides accurate false-alarm rate control even with the relatively small training sample size of 2,000 in this example.
@@
It should be noted that the intended application scenarios in Zhang et al. (2023) involved much larger training sample sizes and/or larger values for the desired $\alpha$. For $\alpha=0.01$ (for example), the method of Zhang et al. (2023) provides much more accurate false-alarm control than in this example with $\alpha=0.001$. With $\alpha=0.001$, the method of Zhang et al. (2023) generally requires that the size of the training subset used to compute the CL is at least 10,000 to accurately estimate the upper 0.001 quantile of the empirical distribution of the $T^{2}$ statistic. In this regard, our bootstrap CL approach can be viewed as an extension of Zhang et al. (2023) to smaller training sample sizes and/or larger desired $\alpha$ values.
@@ 4.2 Nonlinear Oscillator System
To demonstrate our method’s effectiveness in detecting changes in complex nonlinear predictive relationships, we consider a physics-based system that represents a variant of the two-degree-of-freedom nonlinear oscillator with a finite-extensibility coupling spring proposed by Febbo and Machado (2013). The system consists of two masses moving in one dimension, connected by nonlinear springs. Let $p_{1}(t)$ and $p_{2}(t)$ denote the positions at (continuous) time $t$ of the two masses having masses $m_{1}$ and $m_{2}$, respectively, and let $v_{1}(t)=\dot{p}_{1}(t)$ and $v_{2}(t)=\dot{p}_{2}(t)$ denote their velocities. The system evolves according to the following differential equations:
$\ddot{p}_{1}(t)$ $=\frac{1}{m_{1}}\bigl(-k_{1}p_{1}(t)-c_{1}\dot{p}_{1}(t)+k_{3}\phi(p_{1}(t),p_{2}(t))\bigr),$ (4.4)
$\ddot{p}_{2}(t)$ $=\frac{1}{m_{2}}\bigl(-k_{2}p_{2}(t)-c_{2}\dot{p}_{2}(t)-k_{3}\phi(p_{1}(t),p_{2}(t))\bigr),$
where $\phi(p_{1},p_{2})=(p_{1}-p_{2})/(1+|p_{1}-p_{2}|)$ represents the nonlinear coupling force.
To generate the training data, we adopt an ensemble-based sampling scheme. For each observation $i$, the oscillator system in (4.4) is independently initialized with initial conditions $(p_{1}(0),v_{1}(0),p_{2}(0),v_{2}(0))$ drawn i.i.d. from a fixed distribution. The system is then evolved according to (4.4), and the state is recorded at a time $t_{i}$ drawn uniformly from $[0,30]$. This yields an i.i.d. training sample $\{(\mathbf{X}_{i},y_{i})\}_{i=1}^{n}$ with n=3,000n=3{,}000 observations.
The feature vector $\mathbf{X}_{i}\in\mathbb{R}^{4}$ corresponds to the system state at time $t_{i}$ and is given by $\mathbf{X}_{i}=[p_{1}(t_{i}),v_{1}(t_{i}),p_{2}(t_{i}),v_{2}(t_{i})]^{\top}.$ The response variable $y_{i}$ represents the total mechanical energy (kinetic plus potential) of the system at time $t_{i}$, given by
$y_{i}=\frac{1}{2}\bigl(m_{1}v_{1}^{2}(t_{i})+m_{2}v_{2}^{2}(t_{i})\bigr)+\frac{1}{2}\bigl(k_{1}p_{1}^{2}(t_{i})+k_{2}p_{2}^{2}(t_{i})\bigr)+k_{3}\phi\!\left(p_{1}(t_{i}),p_{2}(t_{i})\right)+\varepsilon_{i},$ (4.5)
where $\varepsilon_{i}\sim\mathcal{NID}(0,\sigma^{2})$ represents measurement noise. For the training data, we used the system parameters $m_{1}=1.0$, $m_{2}=2.0$, $k_{1}=1.0$, $k_{2}=2.0$, $k_{3}=1.5$, $c_{1}=0.1$, and $c_{2}=0.2$.
To simulate concept drift, we generate new observations from the same oscillator system but with modified physical parameters $m_{1}^{\prime}=1.1m_{1},m_{2}^{\prime}=1.2m_{2},k_{1}^{\prime}=1.3k_{1}$, while holding the remaining parameters fixed. These changes represent interpretable real-world perturbations, such as increased mass (e.g., due to material accumulation) or stiffened springs (e.g., due to thermal effects), and they alter both the system’s natural frequencies and its energy landscape. As a result, the conditional distribution $\mathbb{P}(Y\mid\mathbf{X})$ is systematically shifted, inducing concept drift that our monitoring framework aims to detect.
To model the nonlinear predictive relationship, we fit a feed-forward multilayer perceptron (MLP) with $L=4$ hidden layers of width $H=32$ and ReLU activations, followed by a linear output layer. That is, the network has the form
$h_{0}(\mathbf{x})=\mathbf{x},\qquad h_{\ell}(\mathbf{x})=\mathrm{ReLU}(W_{\ell}h_{\ell-1}(\mathbf{x})+\mathbf{b}_{\ell})\ \ (\ell=1,\ldots,L),\qquad g_{\gamma}(\boldsymbol{\Theta};\mathbf{x})=\mathbf{w}^{\top}h_{L}(\mathbf{x})+b.$
@@
Following Zhang et al. (2023), we monitor only the parameters of the final linear output layer, denoted by $\boldsymbol{\theta}=(\mathbf{w},b)$, while treating the parameters of the hidden (feature-extraction) layers $\boldsymbol{\Theta}_{\mathrm{feat}}=\{(W_{\ell},\mathbf{b}_{\ell})\}_{\ell=1}^{L}$ as fixed during monitoring. This strategy keeps the dimension of the score vector tractable even for deep architectures, while still capturing changes in the predictive relationship $\mathbb{P}(Y\mid\mathbf{X})$ that propagate through the network (see also Yosinski et al. (2014)). When fitting the model to the training data, all parameters in $\boldsymbol{\Theta}$ are estimated; however, for monitoring, inference is conducted with respect to $\boldsymbol{\theta}$ only, using the penalized Gaussian log-likelihood
@@
$\ell_{\mathrm{pen}}(\boldsymbol{\theta};\mathbf{x}_{i},y_{i})=-\frac{1}{2\sigma^{2}}\bigl(y_{i}-g_{\gamma}(\boldsymbol{\Theta};\mathbf{x}_{i})\bigr)^{2}-\frac{1}{2}\log(2\pi\sigma^{2})-\frac{\gamma}{2n}\|\boldsymbol{\theta}\|^{2},$
with regularization parameter $\gamma=10^{-1}$.
For each observation $(\mathbf{x}_{i},y_{i})$, the score vector is defined as the gradient of the penalized log-likelihood with respect to the monitored parameters $\boldsymbol{\theta}$, evaluated at the fitted values $\hat{\boldsymbol{\theta}}$. Because the network architecture may be deep and highly nonlinear, we do not hand-code this gradient. Instead, the score vector is computed via automatic differentiation by calling loss.backward() in PyTorch. This implementation choice allows the proposed monitoring framework to scale seamlessly to modern deep neural networks without requiring analytic expressions for the score function.
[…]
(a) Low noise ($\sigma=0.03$, 5-fold CV $R^{2}\!\approx\!0.975$)
(b) High noise ($\sigma=0.25$, 5-fold CV $R^{2}\!\approx\!0.137$)